OpenStudy (anonymous):

please help me solve this... 5x+4-3x=2x-2+6

OpenStudy (kc_kennylau):

Infinitely many solutions

OpenStudy (kc_kennylau):

L.H.S. =5x+4-3x =2x+4 R.H.S. =2x-2+6 =2x+4 ∵L.H.S.=R.H.S. ∴The statement is always valid. ∴Infinitely many solutions.

OpenStudy (anonymous):

why thank you kind sir